OpenStudy (anonymous):

You have averaged $57,500.00 for the last five years of work. Your company has agreed to pay you 80 % of the average as your pension. Social security has notified you that you will receive $2100.00 a month. What is your yearly pension?

OpenStudy (amistre64):

its considered bad ettiquette to delete your posts after they have been answered. Just so you know :)

OpenStudy (anonymous):

oh im sorry

OpenStudy (amistre64):

2100 * 12 from SS + 57 500*.8 from the pension i beleive

OpenStudy (amistre64):

2100 * 12+ 57 500*.8 = 71 200 if i see it right

OpenStudy (amistre64):

the reason is that the post belong to the community as a whole to benefit from.

OpenStudy (anonymous):

does * mean times? and ^ divide?

OpenStudy (amistre64):

* means times yes

OpenStudy (anonymous):

i c...i didnt know im sorry

OpenStudy (amistre64):

^ means exponent; / means divide

OpenStudy (anonymous):

oh ok ty

OpenStudy (amistre64):

10/5 = 2

OpenStudy (amistre64):

10^5 = 100 000

OpenStudy (amistre64):

10*5 = 50 lol

OpenStudy (anonymous):

lol ok